OpenStudy (dmezzullo):

Complete the Next Step: A partial solution to the equation 4d − 12 = 12 − 4d is given below. Choose the step that is missing from the solution. 4d − 12 = 12 − 4d 4d − 24 = −4d 3 = d 24 = −8d −24 = 8d −24 = 0d −24 = −8d

OpenStudy (dmezzullo):

@mikala1

OpenStudy (anonymous):

idk sorry

OpenStudy (dmezzullo):

the missing step is btween the 4d and the answer

OpenStudy (dmezzullo):

@RioAnne

OpenStudy (anonymous):

yes that is right

OpenStudy (dmezzullo):

@minnie123

OpenStudy (dmezzullo):

@kymber

OpenStudy (anonymous):

Oh I'm sorry I cant help on this one but I will recommend someone because she's a smart a#$ ... @aussy123

OpenStudy (dmezzullo):

@AriPotta

OpenStudy (aripotta):

d

OpenStudy (anonymous):

Excuse me @minnie123, dont be rude

OpenStudy (anonymous):

Lol tough love, tough love @aussy123

OpenStudy (anonymous):

\[4d-12=12-4d\]\[4d-12-12=-4d\]\[4d-24=-4d\]\[-24=-4d-4d\]\[-24=-8d\]\[3=d\]

OpenStudy (anonymous):

4d − 12 = 12 − 4d 4d − 24 = −4d subtract 4d on both sides -4d -4d ___________________________________________ -24=-8d so It should be D

OpenStudy (anonymous):

I have added some extra steps too..Find the missing one.. :D